# thanks: Corresponding author

### v1

@@ Abstract
In recent years, jumps in the particle circular-orbit radius $r_{\rm c}$ have been proposed as observational signatures of black hole phase transitions, supported by the general consensus that $r_{\rm c}$ increases monotonically with the horizon radius $r_{h}$. However, in this paper, we will show that this monotonicity is not universally valid. By deriving a universal, ensemble-independent criterion for $\mathrm{d}r_{\rm c}/\mathrm{d}r_{h}$ and applying it to the $d$-dimensional Reissner-Nordström-Anti-de Sitter black hole, we reveal a sharp dichotomy: while $r_{\rm c}(r_{h})$ exhibits strictly monotonicity along isobars, it typically becomes non-monotonic along isotherms (except for the photon sphere of Schwarzschild-Anti-de Sitter black hole)–a behavior that has been overlooked in previous studies. Nevertheless, we rigorously prove that within the phase-transition region itself, the monotonicity of $r_{\rm c}(r_{h})$ is preserved. However, the generalizability of this result to other spacetimes requires case-by-case verification, and its validity could have notable implications for the observability of the jump in $r_{\rm c}$. Interestingly, non-monotonicity itself acts as a new observational probe: the monotonicity contrast in $r_{\rm c}(r_{h})$ distinguishes isobaric from isothermal evolution, providing a complementary thermodynamic probe. Furthermore, we show that the critical exponents extracted from $r_{\rm c}$ match those obtained from $r_{h}$, independent of ensemble. Remarkably, all these phenomena are closely tied to the thermodynamic first law.
@@ I Introduction
Black hole thermodynamics has evolved into a central theme of gravitational physics. The Hawking-Page transition first revealed a first-order transition between thermal Anti-de Sitter (AdS) radiation and large AdS black holes [1]. Subsequent work has uncovered thermodynamic phase transitions in various black-hole spacetimes, including Van der Waals-type small/large transitions in AdS black holes [2, 3, 4, 5, 6, 7], thereby establishing a rich thermodynamic phase structure.
Yet direct observation of these thermodynamic phenomena remains limited. In contrast, optical and dynamical features–such as the photon sphere and the innermost stable circular orbit (ISCO)–are directly observable and encode essential geometric information relevant to shadows and accretion physics [8, 9, 10]. The EHT image of M87* made this explicit, revealing an emission ring whose boundary traces the photon orbit [11].
This naturally raises the possibility that black hole thermodynamics could be probed through optical features. Compelling evidence comes from the work of Wei et al., who showed that in charged AdS black holes, both the photon-sphere radius and the minimum impact parameter exhibit Van derWaals-type non-monotonicity and jump discontinuities, behaving as mean-field order parameters [12]. Subsequently, related signatures involving photon orbits and ISCOs have been identified across a variety of spacetimes and modified gravitational theories [13, 14, 15].
@@
A central consideration in these studies is the relationship between the circular-orbit radius $r_{\rm c}$ and the horizon radius $r_{h}$. Across a first-order small/large transition, $r_{h}$ undergoes a discontinuous jumps, and the monotonicity $r_{\rm c}(r_{h})$ ensures that $r_{\rm c}$ also jumps accordingly–making it a natural geometric order parameter (see Appendix A for a brief mathematical derivation). This expectation was later formalized by Zhang et al., who proved monotonic growth of both radii (the photon-sphere radius and ISCO radius) with $r_{h}$ in general static spherical geometries–a result that has been widely applied in subsequent work [16].
@@
However, we find that once more than one thermodynamic parameter is allowed to vary, this monotonicity claim fails. To illustrate this, we present a simple counterexample based on the Schwarzschild-AdS (Sch-AdS) black hole: the photon-sphere radius $r_{\rm ps}=3M$ depends only on the mass $M$, whereas the horizon radius $r_{h}$ depends on both $M$ and $P$, where $P=-\Lambda/(8\pi)$ with $\Lambda$ being the negative cosmological constant. Increasing $P$ decreases $r_{h}$, with $r_{\rm ps}$ unchanged. Consequently, there exist smooth paths in $(M,P)$ parameter space along which $r_{h}$ increases and $r_{\rm ps}$ decreases. Figure 1 provides an explicit example, and Appendix B presents a fully analytic realization. This breakdown of monotonicity in $r_{\rm c}(r_{h})$ has been neglected in previous studies. The reason is straightforward: most analyses have been restricted to the isobaric ensemble, where $P$ is held fixed and only $M$ varies, resulting in both $r_{\rm c}$ and $r_{h}$ typically grow monotonically with $M$. Even in studies that employed the isothermal ensembles (with both $M$ and $P$ varying), the monotonicity of $r_{\rm c}(r_{h})$ was not systematically examined, leading to the omission of non-monotonic behavior. Thus, a genuine understanding of how circular orbits encode thermodynamic information requires analyzing the monotonicity of $r_{\rm c}(r_{h})$ along arbitrary parameter paths–precisely the general framework developed in this work.
Figure 1: Behavior of $M(r_{h},P)$ in the Sch-AdS spacetime. Thin black solid curves denote contours of constant mass $M$. Along the dashed trajectory from A to B, the horizon radius $r_{h}$ increases while both $M$ and the photon-sphere radius $r_{\rm ps}=3M$ decrease, showing that $r_{\rm ps}(r_{h})$ is strictly monotonically decreasing along this path.
@@
Our strategy is as follows. In Sec. II, we review the conditions for photon spheres and ISCOs in a general static, spherically symmetric spacetime, and derive a unified expression for $\mathrm{d}r_{\rm c}/\mathrm{d}r_{h}$ along an arbitrary parameter path. In Sec. III, we analyze the monotonicity of $r_{\rm c}(r_{h})$ for $d$-dimensional Reissner-Nordström-AdS (RN-Ads) black holes, focusing on the isobaric ($T$-$S$) and isothermal ($P$-$V$) thermodynamic paths. In Sec. IV, we show that the critical exponents remain invariant under the reparametrization $r_{h}\!\mapsto r_{\rm c}$, demonstrating that $r_{\rm c}$ defines the same universality class as $r_{h}$. Finally, Sec. V summarizes and discusses our main results.
@@ II.1 Review of circular orbits
where $f(r;\lambda)$ is a smooth function of the radial coordinate $r$, while $\lambda=(\lambda_{1},\ldots,\lambda_{n})$ denotes the set of black-hole parameters, such as mass $M$, charge $q$, and so on. The Lagrangian of a particle moving in the spacetime determined by Eq. (1) reads
@@
$\mathcal{R}(r_{c})=0,\qquad\mathcal{R}^{\prime}(r_{c})=0.$ (6)
Solving Eq. (6) for a massless particle with $\epsilon=0$ gives
@@
where $r_{\rm ps}$ denotes the photon-sphere radius. The condition for an unstable photon sphere further requires
$\mathcal{R}^{\prime\prime}(r_{\rm ps})<0,$ (8)
@@
$\Gamma_{\epsilon}(\lambda)\begin{cases}<0,&\epsilon=0,\\[4.0pt] >0,&\epsilon=1.\end{cases}$ (14)
@@ II.2 Derivation of the Monotonicity Criterion
To investigate the monotonic relation between the circular-orbit radius $r_{c}$ and the event-horizon radius $r_{h}$, we first express them in terms of the previously defined functions $\Phi_{\epsilon}(r;\lambda)$ and $f(r;\lambda)$
@@
$\left.\partial_{r}f\right|_{(r_{h};\lambda)}=2\kappa>0,$ (18)
@@
Based on this, Eq. (15) can in principle be differentiated with respect to $\lambda$. Since the parameter set $\lambda$ generally contains several components, it is convenient to introduce a smooth path $\lambda(t)$ in the parameter space, along which all quantities become functions of a single real variable $t$. This allows us to directly apply the chain rule, yielding
$\frac{\mathrm{d}}{\mathrm{d}t}\,\Phi_{\epsilon}\big(r_{\mathrm{c}}(t);\lambda(t)\big)=\partial_{r_{\mathrm{c}}}\Phi_{\epsilon}\,\dot{r}_{\mathrm{c}}+\langle\nabla_{\lambda}\Phi_{\epsilon},\,\dot{\lambda}\rangle=0,$ (19)
$\frac{\mathrm{d}}{\mathrm{d}t}\,f\big(r_{h}(t);\lambda(t)\big)=\partial_{r_{h}}f\,\dot{r}_{h}+\langle\nabla_{\lambda}f,\,\dot{\lambda}\rangle=0,$ (20)
where the overdot denotes differentiation with respect to $t$, and $\langle\nabla_{\lambda}F,\dot{\lambda}\rangle\equiv\sum_{i=1}^{n}(\partial_{\lambda_{i}}F)\,\dot{\lambda}_{i}$. Combining Eqs. (19) and (20), we finally obtain
@@
In the next section, we apply the criterion (21) to two physically relevant constraints in black hole thermodynamics, namely the isobaric and isothermal processes, and investigate the monotonicity of $r_{\mathrm{c}}(r_{h})$ for the $d$-dimensional RN-AdS black hole.
@@ III Examples of $d$-Dimensional Black Holes under Isobaric and Isothermal Constraints
  • $q=0$: the $d$-dimensional Sch-AdS black hole;
  • $q\neq 0$: the RN-AdS black hole;
@@ III.1 Isobaric Process
The isobaric process corresponds to the $T$-$S$ ensemble commonly employed in black hole thermodynamics. Here, all thermodynamic variables except the mass remain fixed. Thus, we may take $\lambda=(M;\lambda_{\mathrm{fixed}})$ and choose $M$ itself as the path parameter. This leads to $\dot{r}_{h}={\mathrm{d}r_{h}}/{\mathrm{d}M}$. To examine whether $\dot{r}_{h}$ can vanish, we consider the thermodynamic first law
@@
The monotonicity of $r_{\rm c}(r_{h})$ is therefore governed by $\partial_{M}\Phi_{\epsilon}$. Although no fundamental law requires $\partial_{M}\Phi_{\epsilon}\neq 0$, the mass $M$ determines the background geometry and hence controls the response of circular orbits, which generically vary with $M$.
@@
where the explicit expression of $h$ is omitted for brevity. Substituting Eq. (28) into Eq. (27) and using Eqs. (14) and (26), we find
@@
Hence, along an isobaric process, both $r_{\rm ps}(r_{h})$ and $r_{\rm isco}(r_{h})$ increase monotonically. The opposite signs of $\Gamma_{0}$ and $\Gamma_{1}$ are exactly compensated by those of $\partial_{M}\Phi_{0}$ and $\partial_{M}\Phi_{1}$, resulting in a consistent monotonic relation between $r_{\rm c}$ and $r_{h}$. This implies that, during an isobaric process, a small-to-large black hole phase transition necessarily induces a corresponding discontinuous jump in both $r_{\mathrm{ps}}$ and $r_{\mathrm{isco}}$. Such abrupt changes could, in principle, be detected through astronomical observations, providing a potential signature of black hole phase transitions. These findings are consistent with the earlier results reported by Wei, Zhang, et al.
Next, we take $d=4$ as an explicit example and compute the expression of ${\mathrm{d}r_{\rm c}}/{\mathrm{d}r_{h}}$. For the four-dimensional (4D) RN-AdS black hole, one obtains
@@
It is straightforward to verify that the expression of ${\mathrm{d}M}/{\mathrm{d}r_{h}}$ here is consistent with that given in Eq. (25). Using Eq. (27) together with the above results, we find
@@
In the Sch-AdS limit ($q=0$), the above relations reduce to
@@
where $j=1+176\pi P\,r_{\mathrm{isco}}^{2}+1600\pi^{2}P^{2}r_{\mathrm{isco}}^{4}$. Further taking the Schwarzschild limit ($q=0$, $P=0$), one obtains
@@
which are consistent with the well-known relations $r_{\rm ps}=3M$, $r_{\rm isco}=6M$, and $r_{h}=2M$.
In Figs. 2 and 3, we present the numerically computed profiles of $r_{\mathrm{ps}}(r_{h})$ and $r_{\mathrm{isco}}(r_{h})$ for the four-dimensional case. As shown, $r_{\mathrm{c}}(r_{h})$ remains globally monotonic for the Sch-AdS black hole, whereas for the RN-AdS black hole it first decreases and then increases, with the extremum occurring precisely at the point where $\mathrm{d}M/\mathrm{d}r_{h}=\kappa=0$. Truncating the curve at this point, $r_{\mathrm{c}}(r_{h})$ retains its local monotonicity. These numerical results are in complete agreement with Eqs. (33)–(36).
Figure 2: Behavior of $r_{\rm ps}(r_{h})$ along isobaric paths in $d=4$. For the Sch-AdS case, the pressure is likewise set to the value $P=0.8\,P_{\mathrm{c}}$ corresponding to the $q=1$ reference point. The same choice is adopted for subsequent figures and will not be repeated.

### v2

@@ Abstract
Recent studies propose that black hole phase transitions can be encoded in the circular orbit radius of particles. In this paper, we systematically investigate the reliability of this encoding mechanism. We find that this mechanism is highly reliable in the isobaric ensemble, whereas it may break down in the isothermal ensemble. It turns out that the reliability of this mechanism is directly controlled by the first law of black hole thermodynamics. Interestingly, even if this encoding mechanism fails, we prove that, for any black hole exhibiting criticality, the first law can ensure that, near the critical point, the coexistence gap of the circular orbit radius remains a reliable order parameter and yields exactly the same critical exponents as the standard thermodynamic order parameter. Our results provide a potential way to identify the thermodynamic ensemble of a black hole, and reveal a deeper connection between gravitational geometry and thermodynamics.
@@ I Introduction
Black hole thermodynamics provides a unique arena for exploring the interplay among gravity, quantum physics, and statistical mechanics. It began as a formal analogy between the laws of black hole mechanics and ordinary thermodynamics, and acquired genuine physical content through the generalized second law and Hawking radiation [1, 2, 3, 4]. Subsequent developments connected it to broader areas of physics, for example microscopic state counting in string theory and holography in AdS/CFT [5, 6, 7]. Particularly influential progress has come from AdS black holes in the extended phase space, where the cosmological constant is interpreted as a pressure and the mass as an enthalpy. Within this framework, AdS black holes were found to possess a rich phase structure, including Van der Waals-like phase transitions and criticality, as well as more exotic behaviors in modified gravity and nonlinear electrodynamics [8, 9, 10, 11, 12, 13, 14, 15, 16].
These advances further confirm that black hole thermodynamics carries genuine physical content and naturally raise the question how it can be tested. A particularly compelling possibility comes from horizon-scale imaging. The Event Horizon Telescope has imaged the shadows of the supermassive black holes in M87 and at the Galactic center, opening a new observational window on strong-field optical phenomena [17, 18]. Motivated by this breakthrough, recent theoretical work has begun to explore how thermodynamic phase transitions may leave imprints on optical observables. In particular, Wei et al. showed that in the isobaric ensemble the photon sphere radius can serve as an effective probe of the small-/large-black hole phase transition, suggesting that thermodynamic evolution may be reflected in observable image features [19, 20]. Subsequently, related signatures involving photon spheres and innermost stable circular orbits (ISCOs) have been identified across a variety of spacetimes and modified theories of gravity [21, 22, 23, 24].
@@
A key ingredient in these studies is the monotonicity of the mapping between the circular orbit radius $r_{\rm c}$ and the horizon radius $r_{h}$. Across a first-order small/large phase transition, $r_{h}$ undergoes a discontinuous jump, and the monotonic mapping $r_{\rm c}(r_{h})$ ensures that $r_{\rm c}$ also jumps accordingly, thereby making $r_{\rm c}$ a natural geometric order parameter (see Appendix A for a brief mathematical derivation). This expectation was later formalized by Zhang et al., who proved a monotonic increase of the photon sphere and ISCO radii with $r_{h}$ in general static, spherically symmetric geometries, and this result has been widely adopted in subsequent work [25, 26].
@@
However, we find that once more than one thermodynamic parameter is allowed to vary, the claim of monotonic increase may fail. To illustrate this, we present a simple counterexample based on the Schwarzschild–AdS (Sch–AdS) black hole: the photon sphere radius $r_{\rm ps}=3M$ depends only on the mass $M$, whereas the horizon radius $r_{h}$ depends on both $M$ and the pressure $P=-\Lambda/(8\pi)$ associated with the negative cosmological constant $\Lambda$. Increasing $P$ decreases $r_{h}$, with $r_{\rm ps}$ unchanged. Consequently, there exist smooth paths in $(M,P)$ parameter space along which $r_{h}$ increases and $r_{\rm ps}$ decreases. Figure 1 provides an explicit path example, and Appendix B presents an analytic example by constructing an explicit parametric path. This breakdown of monotonic increase in $r_{\rm c}(r_{h})$ has been neglected in previous studies, and the reason is straightforward: Most analyses have been restricted to the isobaric ensemble, where $P$ is held fixed and only $M$ varies, resulting in both $r_{\rm c}$ and $r_{h}$ typically increasing monotonically with $M$. Even when isothermal ensembles were considered (with both $M$ and $P$ varying), the monotonicity of $r_{\rm c}(r_{h})$ was not examined systematically, so possible monotonic decrease or nonmonotonicity was overlooked. Thus, understanding how circular orbits encode black hole phase transitions and other thermodynamic information calls for analyzing the monotonicity of $r_{\rm c}(r_{h})$ along general parameter paths, which is precisely the focus of this work.
@@
Our strategy is as follows. In Sec. II, we review the conditions for the photon sphere and the ISCO in a general static, spherically symmetric spacetime and derive a unified expression for $\mathrm{d}r_{\rm c}/\mathrm{d}r_{h}$ along an arbitrary path in parameter space. In Sec. III, we analyze the monotonicity of $r_{\rm c}(r_{h})$ for $d$-dimensional Reissner-Nordström–AdS (RN–AdS) black holes, focusing on isobaric and isothermal thermodynamic paths. In Sec. IV, we show that the critical exponents are invariant under the reparametrization $r_{h}\!\mapsto r_{\rm c}$, so that $r_{\rm c}$ and $r_{h}$ yield the same universality class. Finally, Sec. V summarizes and discusses our main results. Throughout this work we employ natural units with $c=G=1$.
Figure 1: Behavior of $M(r_{h},P)$ in the Sch–AdS spacetime. Thin black solid curves denote contours of constant mass $M$. Along the dashed trajectory from A to B, the horizon radius $r_{h}$ increases while both $M$ and the photon sphere radius $r_{\rm ps}=3M$ decrease, showing that $r_{\rm ps}(r_{h})$ is strictly monotonically decreasing along this path.
@@ II.1 Review of circular orbits
where $f(r;\lambda)$ is a smooth function of the radial coordinate $r$, while $\lambda=(\lambda_{1},\ldots,\lambda_{n})$ denotes the set of black hole parameters, such as mass $M$, charge $q$, and so on. The Lagrangian of a particle moving in the spacetime determined by Eq. (1) reads
@@
$\mathcal{R}(r_{c})=0,\qquad\mathcal{R}^{\prime}(r_{c})=0,$ (6)
where the prime denotes differentiation with respect to $r$.
Solving Eq. (6) for a massless particle with $\epsilon=0$ gives (see Appendix C for a detailed derivation)
@@
where $r_{\rm ps}$ denotes the photon sphere radius. The condition for an unstable photon sphere further requires
$\mathcal{R}^{\prime\prime}(r_{\rm ps})>0,$ (8)
@@
$\Gamma_{\epsilon}\,\begin{cases}<0,&\epsilon=0,\\[4.0pt] >0,&\epsilon=1.\end{cases}$ (14)
@@ II.2 Derivation of the Monotonicity Criterion
To investigate the monotonic relation between the circular orbit radius $r_{c}$ and the event-horizon radius $r_{h}$, we first express them in terms of the previously defined functions $\Phi_{\epsilon}(r;\lambda)$ and $f(r;\lambda)$
@@
$\left.\partial_{r}f\right|_{(r_{h};\lambda)}=2\kappa\,\neq\,0,$ (18)
@@
Based on this, Eq. (15) can in principle be differentiated with respect to $\lambda$. Since the parameter set $\lambda$ generally contains several components, it is convenient to introduce a smooth path $\lambda(\tau)$ in the parameter space, along which all quantities become functions of a single real variable $\tau$. This allows us to directly apply the chain rule, yielding
$\frac{\mathrm{d}}{\mathrm{d}\tau}\,\Phi_{\epsilon}\big(r_{\mathrm{c}}(\tau);\lambda(\tau)\big)=\partial_{r_{\mathrm{c}}}\Phi_{\epsilon}\,\dot{r}_{\mathrm{c}}+\langle\nabla_{\lambda}\Phi_{\epsilon},\,\dot{\lambda}\rangle=0,$ (19)
$\frac{\mathrm{d}}{\mathrm{d}\tau}\,f\big(r_{h}(\tau);\lambda(\tau)\big)=\partial_{r_{h}}f\,\dot{r}_{h}+\langle\nabla_{\lambda}f,\,\dot{\lambda}\rangle=0,$ (20)
where the overdot denotes differentiation with respect to $\tau$, and $\langle\nabla_{\lambda}F,\dot{\lambda}\rangle\equiv\sum_{i=1}^{n}(\partial_{\lambda_{i}}F)\,\dot{\lambda}_{i}$. Combining Eqs. (19) and (20), we finally obtain
@@
In the next section, we apply the criterion (21) to two physically relevant constraints in black hole thermodynamics, namely the isobaric and isothermal processes, and investigate the monotonicity of $r_{\mathrm{c}}(r_{h})$ for the $d$-dimensional RN–AdS black hole.
@@ III Examples of $d$-Dimensional Black Holes under Isobaric and Isothermal Constraints
  • $q=0$: the $d$-dimensional Sch–AdS black hole;
  • $q\neq 0$: the RN–AdS black hole;
@@ III.1 Isobaric Process
The isobaric process corresponds to the $T$-$S$ ensemble commonly employed in black hole thermodynamics. Here, all thermodynamic variables except the mass remain fixed. Thus, we may take $\lambda=(M;\lambda_{\mathrm{fixed}})$ and choose $M$ itself as the path parameter. This leads to $\dot{r}_{h}={\mathrm{d}r_{h}}/{\mathrm{d}M}$. To examine whether $\dot{r}_{h}$ can vanish, we consider the first law of black hole thermodynamics
@@
The monotonicity of $r_{\rm c}(r_{h})$ is therefore governed by $\partial_{M}\Phi_{\epsilon}$. Although no fundamental law requires $\partial_{M}\Phi_{\epsilon}\neq 0$, the mass $M$ fixes the background geometry and thus controls the circular orbits, so they generically vary with $M$. Based on this consideration and the universality of the first law, we expect that for almost all black holes the mapping $r_{\rm c}(r_{h})$ is monotonic in the isobaric ensemble. As a result, when a phase transition drives the horizon radius to jump from the small-black hole branch $r_{h,s}$ to the large-black hole branch $r_{h,l}$, the circular orbit radius undergoes a corresponding jump from $r_{\rm c}(r_{h,s})$ to $r_{\rm c}(r_{h,l})$. This makes the circular orbit radius a robust and reliable encoder of the phase transition information.
@@
where the explicit expression of $h$ is omitted for brevity. Substituting Eq. (28) into the criterion (27) and using Eqs. (14) and (26), we find
@@
Hence, along an isobaric process, both $r_{\rm ps}(r_{h})$ and $r_{\rm isco}(r_{h})$ increase monotonically. The opposite signs of $\Gamma_{0}$ and $\Gamma_{1}$ are exactly compensated by those of $\partial_{M}\Phi_{0}$ and $\partial_{M}\Phi_{1}$, resulting in a consistent monotonic relation between $r_{\rm c}$ and $r_{h}$. These findings are consistent with the earlier results reported by Wei, Zhang, et al.
Next, we take $d=4$ as an explicit example and compute the expression of ${\mathrm{d}r_{\rm c}}/{\mathrm{d}r_{h}}$. For the four-dimensional RN–AdS black hole, one obtains
@@
It is straightforward to verify that the expression of ${\mathrm{d}M}/{\mathrm{d}r_{h}}$ here is consistent with that given in Eq. (25). Using the criterion (27) together with the above results, we find
@@
In the Sch–AdS limit ($q=0$), the above relations reduce to
@@
where $j=1+176\pi P\,r_{\mathrm{isco}}^{2}+1600\pi^{2}P^{2}r_{\mathrm{isco}}^{4}$. Further, in the limit where both $q$ and $P$ approach zero, one obtains
@@
which are consistent with the well-known relations $r_{\rm ps}=3M$, $r_{\rm isco}=6M$, and $r_{h}=2M$ in the Schwarzschild case.
In Figs. 2 and 3, we present the numerically computed profiles of $r_{\mathrm{ps}}(r_{h})$ and $r_{\mathrm{isco}}(r_{h})$ for the four-dimensional case. In both panels, $P_{c}$ denotes the pressure at the RN–AdS critical point, obtained from the criticality conditions $\partial T/\partial r_{h}=\partial^{2}T/\partial r_{h}^{2}=0$. As illustrated, $r_{\mathrm{c}}(r_{h})$ remains globally monotonic for the Sch–AdS black hole, whereas for the RN–AdS black hole, it first decreases and then increases, with the extremum exactly at the point where $\mathrm{d}M/\mathrm{d}r_{h}=\kappa=0$. Truncating the curve at this point, $r_{\mathrm{c}}(r_{h})$ retains its local monotonicity. These numerical results are in complete agreement with Eqs. (33)–(36).
Figure 2: Behavior of $r_{\rm ps}(r_{h})$ along isobaric paths in $d=4$. For the Sch–AdS case, the pressure is likewise set to the value $P=0.8\,P_{\mathrm{c}}$, where $P_{\mathrm{c}}$ is the critical pressure of the RN–AdS solution with $q=1$. The same choice is adopted for subsequent figures and will not be restated hereafter.